OpenStudy (anonymous):

Write the equations in slope intercept form. x+2y=5 3x-15=-6y

OpenStudy (anonymous):

so we know slope intercept form is y=mx+b so we are try to isolate y so for the first one

OpenStudy (anonymous):

x+2y=5 we subtract x from both sides -x -x 2y=5-x then we divide both sides by 2 and we have \[y=\frac{ 5-x }{ 2 }\]

OpenStudy (anonymous):

now we can make this look more like y=mx+b by splitting up the fraction and putting the x first so we have\[y=-\frac{ 1 }{ 2 }x+\frac{ 5 }{ 2 }\]

OpenStudy (anonymous):

and it is the same process for the second one but for the second one all we have to do is divide by -6

OpenStudy (anonymous):

So on the second one you would have y=1/2X -2/3 ?